# ST-FlowNet: An Efficient Spiking Neural Network for Event-Based Optical Flow Estimation

### v1

@@ Abstract
Spiking Neural Networks (SNNs) have emerged as a promising tool for event-based optical flow estimation tasks due to their ability to leverage spatio-temporal information and low-power capabilities. However, the performance of SNN models is often constrained, limiting their application in real-world scenarios. In this work, we address this gap by proposing a novel neural network architecture, ST-FlowNet, specifically tailored for optical flow estimation from event-based data. The ST-FlowNet architecture integrates ConvGRU modules to facilitate cross-modal feature augmentation and temporal alignment of the predicted optical flow, improving the network’s ability to capture complex motion dynamics. Additionally, to overcome the challenges associated with training SNNs, we introduce a novel approach to derive SNN models from pre-trained artificial neural networks (ANNs) through ANN-to-SNN conversion or our proposed BISNN method. Notably, the BISNN method alleviates the complexities involved in biological parameter selection, further enhancing the robustness of SNNs in optical flow estimation tasks. Extensive evaluations on three benchmark event-based datasets demonstrate that the SNN-based ST-FlowNet model outperforms state-of-the-art methods, delivering superior performance in accurate optical flow estimation across a diverse range of dynamic visual scenes. Furthermore, the inherent energy efficiency of SNN models is highlighted, establishing a compelling advantage for their practical deployment. Overall, our work presents a novel framework for optical flow estimation using SNNs and event-based data, contributing to the advancement of neuromorphic vision applications.
@@ 1 Introduction
As the third generation of neural networks (Maass, 1997), spiking neural networks (SNNs) have garnered increasing attention in recent years (Zhang et al., 2022; Luo et al., 2023; Zheng et al., 2021). Unlike conventional artificial neural networks (ANNs), SNNs employ bio-inspired neurons and discrete spike trains to mimic the complex spatiotemporal dynamics of the human brain (Zheng et al., 2024; Sun et al., 2024). These characteristics enable SNNs to achieve competitive performance across a wide range of neuromorphic intelligent tasks (Ussa et al., 2024; Hagenaars et al., 2021; Yu et al., 2023a). Furthermore, the event-driven information coding and transmission mechanisms ensure lower power consumption, thus enhancing the feasibility of SNNs for hardware implementation (Han and Roy, 2020; Yao et al., 2023).
@@
Optical flow estimation is a fundamental topic in the field of computer vision and has extensive applications (Fleet and Weiss, 2006; Ilg et al., 2017; Vandal and Nemani, 2023), particularly in motion-related intelligent tasks such as object segmentation (Zitnick et al., 2005), video enhancement (Fan et al., 2021; Wang et al., 2020), and object tracking (Kale et al., 2015; Ussa et al., 2024). However, prevalent research has predominantly concentrated on optical flow estimation from frame-based images captured by conventional cameras, resulting in significant performance degradation in challenging scenarios characterized by high-speed motion or unfavorable lighting conditions (Otte and Nagel, 1994; Zhai et al., 2021). Event-based neuromorphic cameras, capable of asynchronously recording changes in light intensity within a high dynamic range of illumination (Han et al., 2020), respond to these challenging scenarios effectively and exhibit potential for minimizing energy consumption, thereby presenting an attractive prospect for deployment on edge devices (Yu et al., 2023a). Furthermore, drawing inspiration from inherent imaging mechanisms, event-based images can circumvent errors introduced by assumptions related to the conservation of pixel intensity in optical flow estimation (Gaur, 2022; Barron et al., 1994). Thus, the advantages of event-based optical flow estimation are significant when compared with conventional approaches (Benosman et al., 2012; Zhu et al., 2018b; Mueggler et al., 2017; Scheerlinck et al., 2020).
@@
Owing to the structural specificity inherent in the event modality, current approaches typically involve the reconstruction of event data into frame-based images for the estimation tasks handled by conventional ANN models (Zhu et al., 2018a; Dosovitskiy et al., 2015; Ilg et al., 2017). Despite some notable progress, these approaches often neglect potential errors inherent in the reconstruction process. Representatively, the resulting reconstructed images frequently exhibit pronounced motion blur in spatial visual cues, particularly as the sampling time window lengthens (Hagenaars et al., 2021; Stoffregen et al., 2020; Tian and Andrade-Cetto, 2022). Concurrently, the valuable dynamic features embedded in the temporal domain remain underutilized. In contrast, an SNN model, comprising spiking neurons as fundamental units, accepts spiking events as input and uses firing spikes as a medium for information propagation and presentation (Ghosh-Dastidar and Adeli, 2009; Tavanaei et al., 2019; Cai et al., 2024). The SNN is an essential tool for event-based feature extraction (Sun et al., 2024; Yu et al., 2022). Consequently, a natural solution to address the aforementioned challenges involves developing an SNN model specifically tailored for event-based optical flow estimation. However, training SNNs using the standard backpropagation (BP) method is hindered by the non-differentiable nature of the spikes (Wu et al., 2018), resulting in constrained performance for real-world tasks.
@@
To fully exploit the potential of event data in optical flow estimation, it is imperative to effectively address two key challenges: (1) developing a novel architecture capable of simultaneous spatio-temporal feature extraction based on event data; and (2) proposing a novel model training strategy for achieving superior optical flow estimation performance. Thus far, limited research has delved into these pertinent issues. Therefore, this work introduces a novel method to effectively tackle these challenges (shown in Fig. 1). We present our ST-FlowNet architecture tailored specifically to estimate event-based optical flow. By incorporating the ConvGRU layers (Ballas et al., 2016), ST-FlowNet achieves cross-scale fusion of dynamic optical flow features from event data in the temporal dimension. In contrast to prior models inspired by pyramidal architectures (Dosovitskiy et al., 2015; Ilg et al., 2017), ST-FlowNet employs a more streamlined decoder structure, enabling direct decoding of latent information across the entire multi-scale feature space. Following training on the ANN model, the SNN model for optical flow estimation can be derived through two strategies: the ANN-to-SNN conversion method (A2S) or our proposed bio-information-fused training strategy (BISNN). Notably, our BISNN approach achieves parameter-free model conversion while preserving the performance of optical flow estimation.
@@ 2 Related Work
Optical flow estimation is a fundamental task in the field of computer vision (Fleet and Weiss, 2006; Ilg et al., 2017), garnering significant research attention. In this section, we initially delve into the evolution of event-based optical flow estimation. Subsequently, our focus shifts to the models built by SNNs, which demonstrate notable proficiency in extracting spatio-temporal features from event-based images.
@@ 2.1 Event-based Optical Flow Estimation
Considering the distinctive attributes of event images in contrast to the RGB modality, early research focused on innovating paradigms for event-based optical flow computation, but only achieved limited success in several simple scenarios (Benosman et al., 2012, 2013). The introduction of large-scale benchmark datasets has facilitated applying deep learning models to event-based optical flow estimation tasks (Zhu et al., 2018b; Mueggler et al., 2017; Scheerlinck et al., 2020), significantly enhancing the efficacy of optical flow estimation and reducing the challenges associated with model design (Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022; Zhu et al., 2018a). Previous research attempted to use encoder-decoder networks to decouple spatio-temporal features embedded in event data across multiple resolutions (Zhu et al., 2018a). A series of variants emerged through the refinement of network structures and adjustments to loss functions, achieving reliable and accurate optical flow estimation facilitated by self-supervised and end-to-end learning methods (Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022). However, considering the high temporal resolution characteristics inherent in event data, models founded on general convolution structures struggle to comprehensively use the temporal features of the event data, thereby limiting the effectiveness of event-based optical flow estimation (Rueckauer and Delbruck, 2016).
@@ 2.2 SNN Models for Optical Flow Estimation
Considering the inherent advantages of SNNs in spatio-temporal feature extraction (Ghosh-Dastidar and Adeli, 2009; Tavanaei et al., 2019), our work focuses on models using spiking neurons which are naturally adept at capturing spatio-temporal visual cues embedded in event data for precise optical flow estimation. Recently, numerous models based on spiking neurons have been proposed for event-based optical flow estimation (Diehl and Cook, 2015; Paredes-Vallés et al., 2019; Orchard et al., 2013; Haessig et al., 2018; Hagenaars et al., 2021; Zhang et al., 2023). Similar to traditional ANN models, shallow convolutional SNN models trained using spike-time-dependent-plasticity learning methods (Diehl and Cook, 2015) have been introduced, demonstrating promising performance for relatively simple tasks (Paredes-Vallés et al., 2019; Orchard et al., 2013; Haessig et al., 2018). To enhance SNN model performance in complex real-world scenarios, a logical approach involves constructing hybrid models that leverage the strengths of both ANNs and SNNs (Diehl and Cook, 2015). Additionally, drawing inspiration from the BP-style direct training method for SNN models (Wu et al., 2018), a series of fully spiking models have also attained state-of-the-art performance on event-based data when compared with advanced ANN models (Hagenaars et al., 2021; Zhang et al., 2023). However, constrained by training approaches and network architectures, we posit that event-based optical flow estimation by SNN models can be further optimized by ANN conversion (Deng and Gu, 2021; Li et al., 2022).
@@ 3 Method
We present an overview of the proposed ST-FlowNet model in Fig. 2. Specifically, ST-FlowNet builds on the foundation of the end-to-end FlowNet model (Dosovitskiy et al., 2015; Ilg et al., 2017), characterized by encoder and decoder layers. To effectively leverage the event modality, which inherently possesses rich spatio-temporal information, we incorporate ConvGRU layers (Cho et al., 2014) for sequential feature argumentation and alignment. Additionally, we use SNN models derived from pre-trained ANN models to enhance the ST-FlowNet’s proficiency in spatio-temporal feature extraction. In this section, we elaborate on the event representation, network architecture, and training methodology used for ST-FlowNet.
@@ 3.2 Network Architecture
As illustrated in Fig. 2, the ST-FlowNet model is structured in a semi-pyramidal encoder-decoder architecture (Dosovitskiy et al., 2015; Ilg et al., 2017). Diverging from prior approaches, we incorporate ConvGRU layers (Cho et al., 2014) to enhance the model’s sequential feature representation. The ConvGRU layer retains its original architecture (shown in Fig. 2(b)). However, the hidden state, which is used to modify information across different time steps, has been redesigned.
@@
Here, $\mathrm{\xi_{R}}$, $\mathrm{\xi_{Z}}$ and $\mathrm{\xi_{I}}$ represent convolutional operations, $\mathrm{\sigma_{1}}$ and $\mathrm{\sigma_{2}}$ are the sigmoid and tanh activation functions. Intuitively, the candidate hidden state $H_{t}$ balances and retains pertinent information from the previous optical flow and current input (8). The reset gate $R_{t}$ governs the degree to which historical information is removed in $H_{t}$. Ultimately, the augmented input $\tilde{I}_{t}$ undergoes an update via a weighted summation of $O_{t-1}$ and $H_{t}$ (5). $O_{-1}$ is established as 00 at the initial time step $t=0$.
@@ 3.3 Model Training
In this work, we use two types of cross-model transformation methodology, A2S conversion (Bu et al., 2022; Deng and Gu, 2021; Rathi and Roy, 2020; Ding et al., 2021; Deng et al., 2020) and bio-information fusion training, to generate an optimal SNN for optical estimation while preserving the original ANN model’s accuracy.
@@ 3.3.1 ANN-to-SNN conversion
where $a^{l}$ is the output in the $l$-th layer, $w^{l}$ signifies the weight between the $(l-1)$-th and $l$-th layers, and $h(\cdot)$ denotes the activation function. In the SNN model, we use the leaky integrate-and-fire model (Abbott, 1999) (LIF) as the spiking neuron for network construction. Without loss of generality, the iterative form of the membrane potential $v^{l}(t)$ of the LIF neuron is described as follows (Rathi and Roy, 2020; Han and Roy, 2020):
@@
The training process for the optical flow estimation network is conducted using the ANN model. Motivated by the limited availability of ground truth optical flow data, we train our optical flow estimation network through a self-supervised approach (Hagenaars et al., 2021). The comprehensive loss function encompasses two fundamental components: a contrast loss $\mathcal{L}_{\mathrm{contrast}}$ and a smoothness loss $\mathcal{L}_{\mathrm{smooth}}$. The contrast loss $\mathcal{L}_{\mathrm{contrast}}$ uses a reformulated contrast maximization proxy loss to gauge the accuracy of optical flow estimation by assessing the motion compensation performance of an image reconstructed from the predicted optical flow (Hagenaars et al., 2021; Gallego et al., 2019, pp. 12280–12289; Mitrokhin et al., 2018; Zhu et al., 2019). The smoothness loss $\mathcal{L}_{\mathrm{smooth}}$ uses Charbonneir smoothness function (Charbonnier et al., 1994; Zhu et al., 2018a, 2019) to regulate the optical flow variation between neighboring pixels. Consequently, the total loss is defined as:

### v2

@@ Abstract
Spiking neural networks (SNNs) have emerged as a promising tool for event-based optical flow estimation tasks due to their capability for spatio-temporal information processing and low-power computation. However, the performance of SNN models is often constrained, limiting their applications in real-world scenarios. To address this challenge, we propose ST-FlowNet, a novel neural network architecture specifically designed for optical flow estimation from event-based data. The ST-FlowNet architecture integrates ConvGRU modules to facilitate cross-modal feature augmentation and temporal alignment of the predicted optical flow, thereby improving the network’s ability to capture complex motion patterns. Additionally, we introduce two strategies for deriving SNN models from pre-trained artificial neural networks (ANNs): a standard ANN-to-SNN conversion pipeline and our proposed BISNN method. Notably, the BISNN method alleviates the complexities involved in selecting biologically inspired parameters, further enhancing the robustness of SNNs for optical flow estimation tasks. Extensive evaluations on three benchmark event-based datasets demonstrate that the SNN-based ST-FlowNet model outperforms state-of-the-art methods, achieving superior accuracy in optical flow estimation across a diverse range of dynamic visual scenes. Furthermore, the energy efficiency of models also underscores the potential of SNNs for practical deployment in energy-constrained environments. Overall, our work presents a novel framework for optical flow estimation using SNNs and event-based data, contributing to the advancement of neuromorphic vision applications.
@@ 1 Introduction
As the third generation of neural networks (Maass, 1997), spiking neural networks (SNNs) have garnered increasing attention in recent years (Zheng et al., 2021). Unlike conventional artificial neural networks (ANNs), SNNs employ bio-inspired neurons and discrete spike trains to mimic the complex spatiotemporal dynamics of the human brain (Zheng et al., 2024; Sun et al., 2024). These characteristics enable SNNs to achieve competitive performance across a wide range of neuromorphic intelligent tasks (Ussa et al., 2024; Yu et al., 2023a). Furthermore, the event-driven information coding and transmission mechanisms ensure lower power consumption, thus enhancing the feasibility of SNNs for hardware implementation (Yao et al., 2023).
@@
Optical flow estimation is a fundamental topic in the field of computer vision and has extensive applications (Ilg et al., 2017; Vandal and Nemani, 2023), particularly in motion-related intelligent tasks. For instance, by clustering motion patterns in different regions, optical flow can assist object segmentation models in separating the foreground and background (Zitnick et al., 2005). The motion vector of the target object is critically important for predicting the search space in object tracking tasks (Ussa et al., 2024). Optical flow also serves as a compensation tool for frame insertion-based image reconstruction and enhancement (Fan et al., 2021; Wang et al., 2020). However, prevalent research has predominantly concentrated on optical flow estimation from frame-based images captured by conventional cameras, resulting in significant performance degradation in challenging scenarios characterized by high-speed motion or unfavorable lighting conditions (Zhai et al., 2021). Event-based neuromorphic cameras, capable of asynchronously recording changes in light intensity within a high dynamic range of illumination (Han et al., 2020), respond to these challenging scenarios effectively and exhibit potential for minimizing energy consumption, thereby presenting an attractive prospect for deployment on edge devices (Yu et al., 2023a). Furthermore, drawing inspiration from inherent imaging mechanisms, event-based images can circumvent errors introduced by assumptions related to the conservation of pixel intensity in optical flow estimation (Gaur, 2022). Thus, the advantages of event-based optical flow estimation are significant when compared with conventional approaches (Zhu et al., 2018b; Mueggler et al., 2017; Scheerlinck et al., 2020).
@@
Owing to the structural specificity inherent in the event modality, current approaches typically involve the reconstruction of event data into frame-based images for the estimation tasks handled by conventional ANN models (Zhu et al., 2018a; Ilg et al., 2017). Despite some notable progress, these approaches often neglect potential errors inherent in the reconstruction process. Representatively, the resulting reconstructed images frequently exhibit pronounced motion blur in spatial visual cues, particularly as the sampling time window lengthens (Stoffregen et al., 2020; Tian and Andrade-Cetto, 2022). Concurrently, the valuable dynamic features embedded in the temporal domain remain underutilized. In contrast, an SNN model, comprising spiking neurons as fundamental units, accepts spiking events as input and uses firing spikes as a medium for information propagation and presentation (Cai et al., 2024). The SNN is an essential tool for event-based feature extraction (Sun et al., 2024). Consequently, a natural solution to address the aforementioned challenges involves developing an SNN model specifically tailored for event-based optical flow estimation.
Unlike conventional ANN models, training SNN models with standard backpropagation (BP) is challenging due to the non-differentiable nature of spike signals (Wu et al., 2018), leading to suboptimal performance in real-world applications. To address this challenge, several methods have been proposed, which can be classified into four main categories: (1) unsupervised learning inspired by biological neuronal plasticity (Diehl and Cook, 2015), (2) indirect training via ANN-to-SNN conversion (A2S) (Deng and Gu, 2021), (3) spatio-temporal backpropagation (STBP) employing surrogate gradient approximation (Wu et al., 2018), and (4) hybrid training strategies (Sun et al., 2024). Among these, STBP has emerged as the most widely adopted approach, enabling SNN training with a procedure similar to that of ANNs while maintaining competitive accuracy. However, STBP-based SNN models typically require extended temporal windows for effective training, resulting in significant computational overhead. Moreover, approximation errors introduced by surrogate gradients further constrain their performance in complex real-world tasks. Consequently, A2S methods have gained prominence as an alternative, allowing the derivation of SNN models from pre-trained ANNs. Compared to the STBP method, A2S strategy substantially reduces training complexity, particularly for intricate task-specific models. Additionally, hybrid strategies integrating diverse training paradigms or neuronal plasticity mechanisms have attracted growing interest, further enhancing SNN training efficiency and model performance.
@@
To fully exploit the potential of event data in optical flow estimation, it is imperative to effectively address two key challenges: (1) developing a novel architecture capable of simultaneous spatio-temporal feature extraction based on event data; and (2) proposing a novel model training strategy for achieving superior optical flow estimation performance. Thus far, limited research has delved into these pertinent issues. Therefore, this work introduces a novel method to effectively tackle these challenges (shown in Fig. 1). We present our ST-FlowNet architecture tailored specifically to estimate event-based optical flow. By incorporating the ConvGRU layers (Ballas et al., 2016), ST-FlowNet achieves cross-scale fusion of dynamic optical flow features from event data in the temporal dimension. In contrast to prior models inspired by pyramidal architectures (Ilg et al., 2017), ST-FlowNet employs a more streamlined decoder structure, enabling direct decoding of latent information across the entire multi-scale feature space. Following training on the ANN model, the SNN model for optical flow estimation can be derived through two strategies: the A2S method or our proposed bio-information-fused training strategy (BISNN). Notably, our BISNN approach achieves parameter-free model conversion while preserving the performance of optical flow estimation.
@@ 2 Related Work
Optical flow estimation is a fundamental task in the field of computer vision (Ilg et al., 2017), garnering significant research attention. In this section, we initially delve into the evolution of event-based optical flow estimation. Subsequently, our focus shifts to the models built by SNNs, which demonstrate notable proficiency in extracting spatio-temporal features from event-based images.
@@ 2.1 Event-based Optical Flow Estimation
Considering the distinctive attributes of event images in contrast to the RGB modality, early research focused on innovating paradigms for event-based optical flow computation, but only achieved limited success in several simple scenarios (Benosman et al., 2013). The introduction of large-scale benchmark datasets has facilitated applying deep learning models to event-based optical flow estimation tasks (Zhu et al., 2018b; Mueggler et al., 2017), significantly enhancing the efficacy of optical flow estimation and reducing the challenges associated with model design (Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022). Previous research attempted to use encoder-decoder networks to decouple spatio-temporal features embedded in event data across multiple resolutions (Zhu et al., 2018a). A series of variants emerged through the refinement of network structures and adjustments to loss functions, achieving reliable and accurate optical flow estimation facilitated by self-supervised and end-to-end learning methods (Scheerlinck et al., 2020; Tian and Andrade-Cetto, 2022). However, considering the high temporal resolution characteristics inherent in event data, models founded on general convolution structures struggle to comprehensively use the temporal features of the event data, thereby limiting the effectiveness of event-based optical flow estimation.
@@ 2.2 SNN Models for Optical Flow Estimation
Considering the inherent advantages of SNNs in spatio-temporal feature extraction (Tavanaei et al., 2019), our work focuses on models using spiking neurons which are naturally adept at capturing spatio-temporal visual cues embedded in event data for precise optical flow estimation. Recently, numerous models based on spiking neurons have been proposed for event-based optical flow estimation (Paredes-Vallés et al., 2019; Hagenaars et al., 2021; Zhang et al., 2023). Similar to traditional ANN models, shallow convolutional SNN models trained using spike-time-dependent-plasticity learning methods (Diehl and Cook, 2015) have been introduced, demonstrating promising performance for relatively simple tasks (Paredes-Vallés et al., 2019). To enhance SNN model performance in complex real-world scenarios, a logical approach involves constructing hybrid models that leverage the strengths of both ANNs and SNNs (Diehl and Cook, 2015). Additionally, drawing inspiration from the BP-style direct training method for SNN models (Wu et al., 2018), a series of fully spiking models have also attained state-of-the-art performance on event-based data when compared with advanced ANN models (Hagenaars et al., 2021; Zhang et al., 2023). However, constrained by training approaches and network architectures, we posit that event-based optical flow estimation by SNN models can be further optimized by ANN conversion (Deng and Gu, 2021).
3 Preliminary
To facilitate a clearer understanding of our work, this section provides a brief overview of SNNs. To date, a variety of spiking neuron models have been developed to emulate the spatio-temporal dynamics of biological neurons. Among these, the leaky integrate-and-fire (LIF) model has become the foundational computational unit in constructing SNNs, offering a balance between biological plausibility and computational efficiency (Abbott, 1999). The membrane potential dynamics of an LIF neuron are typically governed by the following differential equation:
$\tau\frac{dv(t)}{dt}=V_{\text{rest}}-v(t)+I(t),$ (1)
accompanied by the spike generation mechanism defined as:
$s(t)=\begin{cases}1,&\text{if }v(t)\geq\theta,\\ 0,&\text{otherwise},\end{cases}$ (2)
where $v(t)$, $I(t)$, and $s(t)$ denote the membrane potential, input current, and output spike at time $t$, respectively. The parameters $\tau$, $V_{\text{rest}}$, and $\theta$ represent the membrane time constant, resting potential, and firing threshold. A spike is emitted when the membrane potential $v(t)$ reaches or exceeds the threshold $\theta$, after which the membrane potential is reset.
For implementation convenience, the continuous-time differential Eqs (1) and (2) are typically approximated in discrete-time form. The membrane potential update at each time step can be expressed as:
$v(t)=\alpha(t)\cdot\tilde{v}(t-1)+\beta\cdot I(t),$ (3)
where $V_{\text{rest}}$ is assumed to be zero for simplicity. The parameters $\alpha(t)$ and $\beta$ control the decay of the membrane potential and the scaling of the input current, respectively. Here, $\tilde{v}(t-1)$ represents the reset membrane potential at the previous time step, and is defined as:
$\tilde{v}(t-1)=\begin{cases}v(t-1)\cdot(1-s(t-1)),&\text{for hard reset},\\ v(t-1)-s(t-1)\cdot\theta,&\text{for soft reset},\end{cases}$ (4)
where $s(t-1)$ indicates whether a spike was emitted at the previous time step. In the hard reset mechanism, the membrane potential is set to zero following a spike, while in the soft reset case, it is reduced by the threshold value $\theta$.
@@ 4 Method
We present an overview of the proposed ST-FlowNet model in Fig. 2. Specifically, ST-FlowNet builds on the foundation of the end-to-end FlowNet model (Ilg et al., 2017), characterized by encoder and decoder layers. To effectively leverage the event modality, which inherently possesses rich spatio-temporal information, we incorporate ConvGRU layers (Cho et al., 2014) for sequential feature argumentation and alignment. Additionally, we use SNN models derived from pre-trained ANN models to enhance the ST-FlowNet’s proficiency in spatio-temporal feature extraction. In this section, we elaborate on the event representation, network architecture, and training methodology used for ST-FlowNet.
@@ 4.2 Network Architecture
As illustrated in Fig. 2, the ST-FlowNet model is structured in a semi-pyramidal encoder-decoder architecture (Ilg et al., 2017). Diverging from prior approaches, we incorporate ConvGRU layers to enhance the model’s sequential feature representation. The ConvGRU layer retains its original architecture (shown in Fig. 2(b)). However, the hidden state, which is used to modify information across different time steps, has been redesigned.
@@
Here, $\mathrm{\xi_{R}}$, $\mathrm{\xi_{Z}}$ and $\mathrm{\xi_{I}}$ represent convolutional operations, $\mathrm{\sigma_{1}}$ and $\mathrm{\sigma_{2}}$ are the sigmoid and tanh activation functions. Intuitively, the candidate hidden state $H_{t}$ balances and retains pertinent information from the previous optical flow and current input (Eq.12). The reset gate $R_{t}$ governs the degree to which historical information is removed in $H_{t}$. Ultimately, the augmented input $\tilde{I}_{t}$ undergoes an update via a weighted summation of $O_{t-1}$ and $H_{t}$ (Eq.9). $O_{-1}$ is established as 00 at the initial time step $t=0$.
@@ 4.3 Model Training
In this work, we use two types of cross-model transformation methodology, A2S conversion (Bu et al., 2022; Deng and Gu, 2021; Rathi and Roy, 2020) and hybrid bio-information fusion training, to generate an optimal SNN for optical estimation while preserving the original ANN model’s accuracy.
@@ 4.3.1 ANN-to-SNN conversion
where $a^{l}$ is the output in the $l$-th layer, $w^{l}$ signifies the weight between the $(l-1)$-th and $l$-th layers, and $h(\cdot)$ denotes the activation function. In the SNN model, we use the LIF model (Abbott, 1999) as the spiking neuron for network construction. Without loss of generality, the iterative form of the membrane potential $v^{l}(t)$ of the LIF neuron is described as follows (Rathi and Roy, 2020):
@@
The training process for the optical flow estimation network is conducted using the ANN model. Motivated by the limited availability of ground truth optical flow data, we train our optical flow estimation network through a self-supervised approach (Hagenaars et al., 2021). The comprehensive loss function encompasses two fundamental components: a contrast loss $\mathcal{L}_{\mathrm{contrast}}$ and a smoothness loss $\mathcal{L}_{\mathrm{smooth}}$. The contrast loss $\mathcal{L}_{\mathrm{contrast}}$ uses a reformulated contrast maximization proxy loss to gauge the accuracy of optical flow estimation by assessing the motion compensation performance of an image reconstructed from the predicted optical flow (Hagenaars et al., 2021; Mitrokhin et al., 2018; Zhu et al., 2019). The smoothness loss $\mathcal{L}_{\mathrm{smooth}}$ uses Charbonneir smoothness function (Charbonnier et al., 1994; Zhu et al., 2018a, 2019) to regulate the optical flow variation between neighboring pixels. Consequently, the total loss is defined as: